OpenStudy (anonymous):

The length of the hypotenuse of a right triangle is 145 units. The length of one leg of the triangle is 144. Mike wrote the following step to find the length of the unknown leg: Length of the unknown leg = 1452 − 1442 = 21,025 − 20,736 = 289 units Which statement best explains whether Mike's step is correct or incorrect?

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

The length of the hypotenuse of a right triangle is 24 inches. If the length of one leg is 8 inches, what is the approximate length of the other leg?

OpenStudy (anonymous):

thats another question

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

its 22.6 for the second question

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

i really appreciate it

OpenStudy (anonymous):

np

OpenStudy (anonymous):

im not good at math at all lol

OpenStudy (anonymous):

i am :)

OpenStudy (anonymous):

:T

OpenStudy (anonymous):

lol do u need more help

OpenStudy (anonymous):

yesssh i do

OpenStudy (anonymous):

What is the length of the unknown side of the right triangle? A right triangle with hypotenuse x and legs 21 and 20 21 29 400 441

OpenStudy (wolf1728):

(Length of unknown leg)^2 =145^2 - 14^2 (Length of unknown leg)^2 =21,025 - 20736 (Length of unknown leg)^2 =289 Length of unknown leg = Square Root of 289 Length of unknown leg = 17 Therefore, the method is wrong because it did NOT take the square root of 289

OpenStudy (anonymous):

thx man

OpenStudy (wolf1728):

My answer of 17 applies to the first question you asked.

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

the last question u just asked the answer is 29

OpenStudy (anonymous):

User: TASHAWN JACKSON In Course: M/J Grade 8 Pre-Algebra V16 ( 4028) Instructor: Sharon Blackington Exam: 03.07 Module Three Practice Test Warning: There is a checkbox at the bottom of the exam form that you MUST check prior to submitting this exam. Failure to do so may cause your work to be lost. Question 1 (Multiple Choice Worth 1 points) (03.01 HC) The length of the hypotenuse of a right triangle is 145 units. The length of one leg of the triangle is 144. Mike wrote the following step to find the length of the unknown leg: Length of the unknown leg = 1452 − 1442 = 21,025 − 20,736 = 289 units Which statement best explains whether Mike's step is correct or incorrect? It is incorrect because the length of the unknown side is the square root of 289. It is incorrect because the length of the unknown side is the square root of 41,761. It is correct because the length of the unknown side is the difference of the lengths of the sides. It is correct because the length of the unknown side is the difference of the squares of the sides.

OpenStudy (anonymous):

lol my bad :D

OpenStudy (anonymous):

The length of the hypotenuse of a right triangle is 145 units. The length of one leg of the triangle is 144. Mike wrote the following step to find the length of the unknown leg: Length of the unknown leg = 1452 − 1442 = 21,025 − 20,736 = 289 units Which statement best explains whether Mike's step is correct or incorrect? It is incorrect because the length of the unknown side is the square root of 289. It is incorrect because the length of the unknown side is the square root of 41,761. It is correct because the length of the unknown side is the difference of the lengths of the sides. It is correct because the length of the unknown side is the difference of the squares of the sides.

OpenStudy (anonymous):

Look at triangle ABC. Coordinate grid shows negative 5 to positive 5 on the x axis and y axis at intervals of 1. A triangle ABC is shown with A at ordered pair 4, 5, B at ordered pair 2, 2, and C at ordered pair 4, 2. What is the length of side AB of the triangle? 3 5 Square root of 6 Square root of 13

OpenStudy (anonymous):

lol how do i get more people to help me with my math questions.it sees like i need a teacher and all hands on deck lol

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

@dan815 please help

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@chancemorris123 pls help

OpenStudy (anonymous):

:T

OpenStudy (lynfran):